OpenStudy (anonymous):

A system of equations is shown below. -3x+7y=-16 -9x +5y=16 Part A: Create an equivalent system of equations by replacing one question with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent systems has the same solution as the original systems of equations.

OpenStudy (anonymous):

Please need help ASAP

OpenStudy (math&ing001):

-3*(-3x+7y=-16 ) + (-9x+5y=16) -------------------- -16y= 4*16

OpenStudy (anonymous):

@math&ing001 can you help me with part b?

OpenStudy (anonymous):

@Loser66 ?

OpenStudy (math&ing001):

The 2 systems have the same solution set because you can find the initial equations from those obtained later.

OpenStudy (anonymous):

Oh, is that part B? @math&ing001

OpenStudy (math&ing001):

Yes !

OpenStudy (anonymous):

Thank you!

OpenStudy (math&ing001):

You're welcome :)